Consider the sequence that begins 40, 20, 10, 5

a. based on the information, can this sequence by arithmetic? Geometric? Why?

b. Assume this is a geometric sequence on a graph up to n=6

c. Will the values of the sequence ever become zero or negative? Explain.
Answer #1


40, 20, 1015 a) Arithmetic proogrension An At (m.tld 20- 40 +62-1) d 1 d: 20 Ay= At (m.1) d 10= 40 +(212 od - -30 I da 15 I NRatio of two cousecutive terms is constant - and that is why this is a o n o o - Greauetric progression and after (b) 1 m= 6]the value of an will be decreasing as increment in the value of n. because an would be decreasing, and Overall de creament wo

Please upvote and mention your doubts in the comments if any.